### v1

@@ 1 Introduction
On the positive side, $k$-DDP retains some degree of tractability when the class of digraphs is restricted. For instance, Fortune et al. [12] showed that it is XP with parameter $k$ in DAGs, and this result was later extended to digraphs of bounded directed treewidth by Johnson et al. [14], more precisely by presenting an XP algorithm with both $k$ and the directed treewidth as parameters.
@@
A bramble $\mathcal{B}$ in a digraph $D$ is a collection of strongly connected subgraphs of $D$, called the bags, such that for any $B,B^{\prime}\in\mathcal{B}$ the digraph induced by $V(B)\cup V(B^{\prime})$ is also strong. The order of $\mathcal{B}$ is minimum size of a vertex set of $D$ intersecting all its elements. For an integer $c\geq 1$, we say that $\mathcal{B}$ has congestion $c$ if every vertex of $D$ appears in at most $c$ bags of $\mathcal{B}$. Edwards et al. [10] showed how to construct, in polynomial time, a bramble of order $k$ and congestion $2$ from a bramble of order $f(k)$, albeit with a very large exponential dependency on $k$. This tool is a fundamental part of the main contribution of their paper, where they show that every $(36k^{3}+2k)$-strong digraph containing a bramble of congestion $2$ and size $188k^{3}$ is $(k,2)$-linked. Notice that since a vertex can only hit $c$ elements of a bramble with congestion $c$, the order and size of each such bramble is linearly related to its size. They also show how to find the desired set of paths in polynomial time assuming that the bags of the bramble are given. Since one can always find a bramble of order $t$ in digraphs of directed treewidth at least $13t-$ 1 (see [5] and [2, Theorem 9.4.4]), they conclude that there is a function $g(k)$ such that every $(36k^{3}+2k)$-strong digraph of directed treewidth at least $g(k)$ is $(k,2)$-linked. Finally, as the directed treewidth of a $t$-strong digraph $D$ is at least $t$, they conclude that there is a function $L(k)$ such that every $L(k)$-strong digraph is $(k,2)$-linked.
@@
Our second contribution considers the result by Masařík et al. [19], where it is shown that there is an integer $t_{k}=\mathcal{O}(k^{48}\log^{13}k)$ such that every digraph of directed treewidth at least $t_{k}$ contains a bramble of size at least $k$ and congestion at most $8$. We show how to obtain a polynomial-time algorithm from their proof (assuming that a path system, as defined in section 2, is given) by applying two local changes to it, since many parts of the proof are readily translatable into polynomial-time algorithms.
@@
The local changes that we apply are about finding clique minors in graphs of large average degree, as done by Thomasson [28] and Kostochka [16] (cf. [19, Lemma 2.4]), and an application of the Lovász Local Lemma (henceforth abbreviated as LLL). To find the desired clique minor in polynomial time, it suffices to replace the result used in [19] by the algorithmic construction by Dujmović [9, Theorem 1.1]. The application of the LLL requires more effort, and in what follows we sketch the main ingredients of our approach.
@@
The proof of Masařík et al. [19] follows a series of steps refining a previously obtained object into another one, starting from a path system (see 2.6). At many stages of the proof, the intersection graph of pairs of sets of linkages is tested for its degeneracy. In case it is sufficiently large as a function of $k$, the desired bramble is found through an application of the results of [28, 16] (or, in our case, the algorithm of [9]). Otherwise, more subtle arguments are needed to distinguish between sparse cases. In any case, the construction often relies on a classical application of the LLL to find multicoloured independent sets in graphs that have bounded degeneracy. The original proof of the LLL by Lovász and Erdős [17], however, is not constructive.
Although many advances were made in designing algorithms for the LLL, those results rely on restrictions on the instance that are not guaranteed to hold in the application of [19]. Thus, in order to obtain a polynomial-time algorithm from their proof, in Theorem 4.3 we apply an algorithmic version of the LLL by Chandrasekaran et al. [6] to prove an algorithmic version of a result by Reed and Wood [25] appearing as [19, Lemma 2.5].
@@
The usage of the algorithmic version of the LLL by Chandrasekaran et al. [6] has an impact on how large the directed treewidth must be to ensure that the bramble can be found. We show that for every $\alpha\geq 1$ there is an integer $t^{\prime}_{k}=\mathcal{O}\left((k\cdot\log k^{2})^{\textsf{poly}(\alpha)}\right)$ such that every digraph with directed treewidth at least $t^{\prime}_{k}$ has a bramble of congestion $8$ and size $k$. The choice of $\alpha$ has a large impact in the running time of the algorithm: $\alpha$ is directly proportional to $t^{\prime}_{k}$ and, up to a certain point, the LLL dependent algorithm runs faster as it grows (see the discussion at the end of subsection 4.4). We remark that our contribution does not require any major changes to the proof by Masařík et al. [19]: we follow the exact same steps of their proof, but adjusting the bounds appropriately for applying our result derived from an algorithmic version of the LLL.
@@
It is open if it is possible to improve the $\mathcal{O}(k^{48}\log^{13}k)$ bound from [19] to obtain brambles of size $k$ and fixed congestion, and whether it is possible to obtain brambles of smaller congestion if a larger dependence on $k$ is allowed. An improvement on the bound from [19] implies an improvement to $f(k,\alpha)$ defined above. If a large bramble of even congestion $c$ can be found, then Theorem 2.14 can be applied without any quality loss on the obtained paths. If a bramble of congestion at most $6$ is obtainable, then again Theorem 2.14 can be applied to conclude $(k,c^{\prime})$-linkedness of $k$-strong digraphs with $c^{\prime}\leq 6$.
@@ 2.3 Brambles, linkages, and path systems
Originally, the result of Campos et al. [5] builds a haven of order $k$ in digraphs of directed treewidth at least $k$. Since one can easily construct a bramble of order $\lfloor k/2\rfloor$ from a haven of order $k$ and generate a haven of order $k$ from a bramble of order $k$ (see [21, Chapter 6] for example), we refrain from defining havens and state the next proposition with relation to brambles only, noting that the original statement refers to havens. We also do not need the definition of directed treewidth and thus it is also skipped. The aforementioned XP algorithm [14] was improved to an FPT algorithm, with the same approximation ratio, by Campos et al. [5]. We remark that a sketch of a proof of an FPT algorithm is also available in [2, Theorem 9.4.4], but with a slightly worse approximation ratio.
@@ 2.4 Constant congestion brambles and $(k,c)$-linked digraphs
The proof of 2.8 skips directed treewidth and assumes that an $(a,b)$-path system is given, where both $a$ and $b$ depend polynomially on $k$ (see 2.7). In section 4 we show how to adapt their construction into a polynomial-time algorithm that finds a bramble of congestion $8$ and size at least $k$, assuming that an $(a,b)$-path system is given with $a$ and $b$ also depending polynomially on $k$, but with a worse bound than that of 2.8. This loss is mainly due to an application of a polynomial-time and deterministic algorithm for the Lovász Local Lemma. More precisely, we obtain the following.
@@
The values of $t$ and $\ell$ will be precised latter. A statement in the same shape of 2.8 is given in Theorem 2.15. The requirement of Theorem 2.9 is not much different than the requirement of 2.8 since in their proof the authors use the directed treewidth only to obtain a path system.
@@
such that every $k$-strong digraph $D$ with ${\sf dtw}(D)\geq t_{k}$ is $(k,8)$-linked. Additionally, any given instance $(D,S,T)$ of $(k,8)$-DDP can be solved in time $\textsf{poly}(k,\alpha)\cdot n^{2}$.
Finally, applying 2.10 and Theorem 2.15 we prove the following.
@@
For all $\alpha>1$ and integers $k\geq 1$ there is a polynomial function $g(k,\alpha)$ such that every $g(k,\alpha)$-strong digraph is $(k,8)$-linked.
@@ 4.2 Finding multicoloured independent sets
The following definitions are needed to state [6, Theorem 5]. Let $\mathcal{P}=\{P_{1}\,\ldots,P_{n}\}$ be a collection of mutually independent discrete random variables and, for $i\in[n]$, let $D_{i}$ be the finite domain of $P_{i}$. Let $\mathcal{A}=\{A_{1},\ldots,A_{m}\}$ be a collection of events each of which is determined by a subset of $\mathcal{X}$. The variable set of an event $A\in\mathcal{A}$ is the unique minimal $\mathcal{S}\subseteq\mathcal{P}$ that determines $A$ and is denoted by $\textsf{vbl}(A)$. We denote by $\Gamma(A)$ the set $\{B\in\mathcal{A}\mid\textsf{vbl}(A)\cap\textsf{vbl}(B)\neq\emptyset\}$. Thus, $\Gamma(A)$ denotes the set of events of $\mathcal{A}$ which are not independent from $A$.
@@
We now use 4.2 to derive an algorithmic version of a result by Reed and Wood [25, Lemma 4.3].
@@
By the assumption that each $G[V_{i}\cup V_{j}]$ is $b$-degenerate and since every $b$-degenerate graph with $\ell$ vertices has at most $b\cdot\ell$ edges, we conclude that, for each $i\in[r]$, the number of edges with exactly one endpoint in $V_{i}$ is at most $(r-1)b\cdot 2t$. This, in turn, implies that $|\Gamma(A_{u,v})|\leq 4b(r-1)$ since, by definition, $A_{x,y}\in\Gamma(A_{u,v})$ if $\{x,y\}\cap\{u,v\}\neq\emptyset$ and $\{x,y\}\neq\{u,v\}$. The intuition here is that, for $u\in V_{i}$ and $v\in V_{j}$ with $i\neq j$, the event $A_{u,v}$ is independent from every other event in $\mathcal{A}$ associated with an edge $e\in F$ with endpoints in sets of $\mathcal{V}$ distinct from both $V_{i}$ and $V_{j}$, and thus the degeneracy assumption implies that the bad events are reasonably independent from each other. Defining $d=4b(r-1)t$, we are ready to invoke 4.2.
Since every $\textsf{Pr}(A_{u,v})=1/t^{2}$, to apply 4.2 it suffices to choose $x(A_{u,v})=1/d$ for all $(A_{u,v}\in\mathcal{A})$. Now, we want that
@@
and thus the choice of $t$ guarantees that the conditions of 4.2 are satisfied and a good evaluation is found in polynomial time.
@@ 4.3 The algorithmic construction
Let $G$ be a graph with one vertex for each linkage $L_{i,j}$ of the path system $\mathcal{(P,L,A)}$ with $i\neq j$ and no edges. Formally, $V=\{(i,j)\in[a]\times[a]\mid i\neq j\}$ and $G=(V,\emptyset)$. The authors of [19] apply [19, Lemma 3.5] to $\mathcal{(P,L,A)}$ and thus associate each vertex of $G$ with a threaded linkage of size $b$, as defined in [19]. The goal now is to apply [19, Lemma 3.6] to each of those objects to associate with each vertex of $G$ either a set of closed walks with additional properties, including small congestion, or with an untangled threaded linkage. For that, we need that $b\geq x\cdot d+(d-1)$, for some choice of $x$ and $d$, in order to conclude that the obtained sets of closed walks have size $d$ and the untangled threaded linkages have size at least $x$. Let $\alpha>1$. Choosing
@@
Thus, the lemma is applicable. The set of vertices of $G$ for which the first outcome is obtained, namely the closed walks, is named $Z$. As a consequence of applying [19, Lemma 3.6] to each $(i,j)\in Z$, we can also associate each such vertex with a threaded linkage satisfying some additional properties. Whether each $(i,j)\in Z$ is used to identify one of such threaded linkages or a closed walk depends on the application. For now,
@@
We remark that, so far, we have applied [19, Lemmas 3.5 and 3.6] a total of $a(a-1)$ times each.
They now build two undirected graphs on top of $V$, each depending on the choices of $d_{1}$ and $d_{2}$ with $d_{1}\geq d_{2}$. These choices are done in order to satisfy the LLL conditions, and thus need to be adapted to satisfy the polynomial LLL conditions. Start with $E_{1}=E_{2}=\emptyset$. For each pair $(i,j),(i^{\prime},j^{\prime})\in V$ and $i\in\{1,2\}$, add the pair to $E_{i}$ if ${\sf Int}(L^{\prime}_{i,j}\cup L^{\prime}_{i^{\prime},j^{\prime}})$ is not $d_{i}$-degenerate and define $H_{i}=(V,E_{i})$. Finally, let $M_{1}$ be a maximum matching in $H_{1}\setminus Z$ and $M_{2}$ be a maximum matching in $H_{2}$ excluding every edge with both endpoints in $Z$ plus vertices used by $M_{1}$. For $i\in\{1,2\}$ let $V(M_{i})$ be the set of vertices used by the edges of $M_{i}$. It is shown [19, Claim 4.1] that at least one of these three cases occurs:
@@
The algorithmic version of the sparse winning scenario is obtained by simply following the original proof and applying 4.1 instead of [19, cf. Lemma 2.4].
Lemma 4.5 (sparse winning scenario [19, Lemma 3.2], algorithmic version).
@@
Similarly, the algorithmic version of the sparse winning scenario (wrapped) is obtained by following the original proof and adjusting the bounds to apply Theorem 4.3 instead of [19, Lemma 2.5].
Lemma 4.6 (sparse winning scenario, wrapped [19, Lemma 3.3], algorithmic version).
Let $c$ be the constant from 4.5. Let $D$ be a digraph, let $a,b\geq 1$ be integers and $(P,\mathcal{L},\mathcal{A})$ be an $(a,b)$-path system in $D$. Let $\mathcal{I}$ be a subset of $[a]\times[a]\setminus\{(i,i)\mid i\in[a]\}$ with size at least $0.6a(a-1)$, and assume that there is an integer $d$ such that, for every distinct pair $(i,j),(i^{\prime},j^{\prime})\in\mathcal{I}$ the intersection graph ${\sf Int}(L_{i,j}\cup L_{i^{\prime},j^{\prime}})$ is $d$-degenerate. If $b^{1-\varepsilon}\geq(e\cdot 4a^{2}\cdot d)^{1+\varepsilon}$, then in polynomial time we can find in $D$ a bramble of congestion at most $4$ and size at least
@@
In addition, for case 2. another computation step is needed in the form of the Bowtie Lemma [19, Lemma 3.7], whose proof also naturally yields a polynomial-time algorithm, and the dense winning scenario. Again, the algorithmic version follows by incorporating 4.1 to the original proof.
Lemma 4.7 (dense winning scenario [19, Lemma 3.1], algorithmic version).
@@
For case 1., let $\mathcal{I}=V\setminus(V(M_{1})\cup Z)$. The goal is to apply 4.6 to a path system built from linkages associated with pairs $(i,j)\in\mathcal{I}$. Since the pairs in $Z$ are excluded, every linkage in $\{L^{\prime}_{i,j}\mid(i,j)\in\mathcal{I}\}$ has size at least $(e\cdot 4a^{2}\cdot d_{1})^{\alpha}+1$. Thus we need that

### v2

@@ 1 Introduction
On the positive side, $k$-DDP retains some degree of tractability for restricted classes of digraphs. For instance, Fortune et al. [12] showed that it is XP with parameter $k$ in DAGs, which was later extended to digraphs of bounded directed treewidth by Johnson et al. [15], by presenting an XP algorithm with both $k$ and the directed treewidth as parameters.
@@
A bramble $\mathcal{B}$ in a digraph $D$ is a collection of strongly connected subgraphs of $D$, called bags, such that for any $B,B^{\prime}\in\mathcal{B}$ the digraph induced by $V(B)\cup V(B^{\prime})$ is also strong. The order of $\mathcal{B}$ is the minimum size of a vertex set of $D$ intersecting all its elements. For an integer $c\geq 1$, we say that $\mathcal{B}$ has congestion $c$ if every vertex of $D$ appears in at most $c$ bags of $\mathcal{B}$. Edwards et al. [10] showed how to construct, in polynomial time, a bramble of order $k$ and congestion $2$ from a bramble of order $f(k)$, albeit with a very large exponential dependency on $k$. This tool is a fundamental part of the main contribution of their paper, where they show that every $(36k^{3}+2k)$-strong digraph containing a bramble of congestion $2$ and size $188k^{3}$ is $(k,2)$-linked. Notice that since a vertex can only hit $c$ elements of a bramble with congestion $c$, the order and size of each such bramble is linearly related to its size. They also show how to find the desired set of paths in polynomial time assuming that the bags of the bramble are given. Since one can always find a bramble of order $t$ in digraphs of directed treewidth at least $13t-$ 1 (see [5] and [2, Theorem 9.4.4]), they conclude that there is a function $g(k)$ such that every $(36k^{3}+2k)$-strong digraph of directed treewidth at least $g(k)$ is $(k,2)$-linked. Finally, as the directed treewidth of a $t$-strong digraph $D$ is at least $t$, they conclude that there is a function $L(k)$ such that every $L(k)$-strong digraph is $(k,2)$-linked.
@@
Our second and main contribution considers the result by Masařík et al. [20], where it is shown that there is an integer $t_{k}=\mathcal{O}(k^{48}\log^{13}k)$ such that every digraph of directed treewidth at least $t_{k}$ contains a bramble of size at least $k$ and congestion at most $8$. We show how to obtain a polynomial-time algorithm from their proof (assuming that a path system, as defined in section 2, is given) by applying two local changes to it, since many parts of the proof are readily translatable into polynomial-time algorithms.
@@
The local changes that we apply are about finding clique minors in graphs of large average degree, as done by Thomasson [29] and Kostochka [17] (cf. [20, Lemma 2.4]), and an application of the Lovász Local Lemma, henceforth abbreviated as LLL. To find the desired clique minor in polynomial time, it suffices to replace the result used in [20] by the algorithmic construction by Dujmović [9, Theorem 1.1].
@@
In what follows we sketch the main ingredients of our approach. The proof of Masařík et al. [20] follows a series of steps refining a previously obtained object into another one, starting from a path system (see 2.6). At many stages of the proof, the intersection graph of pairs of sets of linkages is tested for its degeneracy. In case it is sufficiently large as a function of $k$, the desired bramble is found through an application of the results of [29, 17] (or, in our case, the algorithm of [9]). Otherwise, more subtle arguments are needed to distinguish between sparse cases. In any case, the construction often relies on a classical application of the LLL to find multicoloured independent sets in graphs that have bounded degeneracy. The original proof of the LLL by Lovász and Erdős [18], however, is not constructive. From this point forward, we have two options to obtain an algorithmic version of a result by Reed and Wood [26] appearing as [20, Lemma 2.5]. In both cases, the running time of the algorithm is split as the sum of two parts, one of which depends only on the chosen approach to deal with the application of the LLL which, in turn, is applied on a graph whose size depends only on $k$ and on the choice of a constant $\alpha\geq 1$. We remark that this dependence is likely to be the dominant one in the running time of the obtained polynomial-time algorithm.
For the first, we can apply a result by Harris [13, Theorem 1.5] to adapt the proof of [20] into a polynomial-time algorithm while maintaining the same asymptotic bound for the directed treewidth of the input digraph [^1] This approach was suggested to us by David G. Harris after seeing the first version of this article on arXiv, containing only the second approach mentioned here. We would like to thank him for pointing out the existence of 4.2 and for the helpful comments on how to apply it to obtain Theorem 2.16.. The downside of this approach is that, according to Harris [14], in this case a precise bound on the running time of the LLL-dependent part of the algorithm cannot be given without significant technical work. Thus, for the second option we apply an algorithmic version of the LLL by Chandrasekaran et al. [6], which requires more work, to obtain an upper bound on the running time of the LLL-dependent part of the obtained algorithm.
@@
The usage of the algorithmic version of the LLL by Chandrasekaran et al. [6] has an impact on how large the directed treewidth must be to ensure that the bramble can be found. We show that for every $\alpha>1$ there is an integer $t^{\prime}_{k}=\mathcal{O}\left((k\cdot\log k^{2})^{\textsf{poly}(\alpha)}\right)$ such that every digraph with directed treewidth at least $t^{\prime}_{k}$ has a bramble of congestion $8$ and size $k$. The choice of $\alpha$ has a large impact in the running time of the algorithm: $\alpha$ is directly proportional to $t^{\prime}_{k}$ and, up to a certain point, the LLL-dependent algorithm runs faster as it grows (see the discussion at the end of subsection 4.4). We remark that our contribution does not require any major changes to the proof by Masařík et al. [20]: we follow the exact same steps of their proof, but adjusting the bounds appropriately for applying our result derived from an algorithmic version of the LLL.
@@
It is open whether it is possible to improve the $\mathcal{O}(k^{48}\log^{13}k)$ bound from [20] to obtain brambles of size $k$ and fixed congestion, and whether it is possible to obtain brambles of smaller congestion if a larger dependence on $k$ is allowed. An improvement on the bound from [20] implies an improvement to $f(k,\alpha)$ defined above. If a large bramble of even congestion $c$ can be found, then Theorem 2.14 can be applied without any quality loss on the obtained paths. If a bramble of congestion at most $6$ is obtainable, then again Theorem 2.14 can be applied to conclude $(k,c^{\prime})$-linkedness of $k$-strong digraphs with $c^{\prime}\leq 6$.
@@ 2.3 Brambles, linkages, and path systems
Originally, the result of Campos et al. [5] builds a haven of order $k$ in digraphs of directed treewidth at least $k$. Since one can easily construct a bramble of order $\lfloor k/2\rfloor$ from a haven of order $k$ and generate a haven of order $k$ from a bramble of order $k$ (see [22, Chapter 6] for example), we refrain from defining havens and state the next proposition with relation to brambles only, noting that the original statement refers to havens. We also do not need the definition of directed treewidth and thus it is also skipped. The aforementioned XP algorithm [15] was improved to an FPT algorithm, with the same approximation ratio, by Campos et al. [5]. We remark that a sketch of a proof of an FPT algorithm is also available in [2, Theorem 9.4.4], but with a slightly worse approximation ratio.
@@ 2.4 Constant congestion brambles and $(k,c)$-linked digraphs
The proof of 2.8 skips directed treewidth and assumes that an $(a,b)$-path system is given, where both $a$ and $b$ depend polynomially on $k$ (see 2.7). 2.8 can be adapted into a polynomial-time algorithm by applying the result by Harris [13, Theorem 1.5] (cf. 4.2) without any loss on the asymptotic bound of $t_{k}$. However, according to the author [14], significant technical work is needed to precise the running time of the algorithm.
To provide a version of 2.8 with a bound on the running time, in section 4 we show how to adapt their construction into a polynomial-time algorithm that finds a bramble of congestion $8$ and size at least $k$, assuming that an $(a,b)$-path system is given with $a$ and $b$ also depending polynomially on $k$, but with a worse bound than that of 2.8. This loss is mainly due to an application of a polynomial-time and deterministic algorithm for the Lovász Local Lemma. More precisely, we obtain the following.
@@
The values of $t$ and $\ell$ will be precised later. A statement in the same shape of 2.8 is given in Theorem 2.15. The requirement of Theorem 2.9 is not much different than the requirement of 2.8 since in their proof the authors use the directed treewidth only to obtain a path system.
@@
such that every $k$-strong digraph $D$ with ${\sf dtw}(D)\geq t_{k}$ is $(k,8)$-linked. Additionally, any given instance $(D,S,T)$ of $(k,8)$-DDP can be solved in time $\textsf{poly}(k,\alpha)\cdot n^{2}$, where the degree of the polynomial depending on $k$ and $\alpha$ comes from Theorem 4.4.
The running time of Theorem 2.15 depends on Theorem 4.4 and is further discussed in the end of subsection 4.3. Alternatively, using 4.2 we obtain a better bound on $t_{k}$ (the same as Theorem 2.15 with $\alpha=1$) but the degree of $\textsf{poly}(k,\alpha)$ is unknown.
Theorem 2.16.
For all $\alpha>1$ and integer $k\geq 1$ there is an integer
$t_{k}=\mathcal{O}\left(k^{96}\cdot(\log k^{2})^{13}\right)$
such that every $k$-strong digraph $D$ with ${\sf dtw}(D)\geq t_{k}$ is $(k,8)$-linked. Additionally, any given instance $(D,S,T)$ of $(k,8)$-DDP can be solved in polynomial time.
Finally, applying 2.10, Theorem 2.15, Theorem 2.16 we prove the following.
@@
For all $\alpha\geq 1$ and integers $k\geq 1$ there is a polynomial function $g(k,\alpha)$ such that every $g(k,\alpha)$-strong digraph is $(k,8)$-linked.
@@ 4.2 Finding multicoloured independent sets
As mentioned in the previous sections, we have two options to find the multicolored independent sets in our setup. We can apply the following result by Harris [13, Theorem 1.5] to obtain Theorem 2.16, which has a better bound for $t_{k}$ with the downside of not providing a precise bound of the running time.
Proposition 4.2 (Harris [13]).
Let $G$ be a digraph with maximum degree $\Delta$ and $\mathcal{V}$ be a partition of $V(G)$ into sets $V_{1},\ldots,V_{r}$ of size exactly $t$. Let $\varepsilon>0$ be an arbitrary constant and suppose that $t\geq(2+\varepsilon)\Delta$. Then there is a polynomial-time algorithm that finds in $G$ an independent set $\{v_{1},\ldots,v_{r}\}$ such that $v_{i}\in V_{i}$ for each $i\in[r]$.
It is easy to apply 4.2 under the assumption that $G[V_{i}\cup V_{j}]$ is $b$-degenerate for any distinct $i,j\in[r]$. It suffices to notice that, since there are at most $2t(r-1)b$ edges leaving each $V_{i}$ in this case, at least $t/2$ vertices of each such set have maximum degree at most $4t(r-1)b$. Thus one can apply 4.2, for example, under the assumption that $t\geq 17(r-1)b$. Since this bound on $t$ is within a constant factor of the original conditions needed to apply [20, Lemma 2.5] (which asks that $t\geq 4e(r-1)b$), to obtain Theorem 2.15 it suffices to analyse the running time of the remaining steps of the original construction, as done in the end of subsection 4.3.
To obtain Theorem 2.15, additional work is need to apply [6, Theorem 5]. We need the following definitions. Let $\mathcal{P}=\{P_{1}\,\ldots,P_{n}\}$ be a collection of mutually independent discrete random variables and, for $i\in[n]$, let $D_{i}$ be the finite domain of $P_{i}$. Let $\mathcal{A}=\{A_{1},\ldots,A_{m}\}$ be a collection of events each of which is determined by a subset of $\mathcal{X}$. The variable set of an event $A\in\mathcal{A}$ is the unique minimal $\mathcal{S}\subseteq\mathcal{P}$ that determines $A$ and is denoted by $\textsf{vbl}(A)$. We denote by $\Gamma(A)$ the set $\{B\in\mathcal{A}\mid\textsf{vbl}(A)\cap\textsf{vbl}(B)\neq\emptyset\}$. Thus, $\Gamma(A)$ denotes the set of events of $\mathcal{A}$ which are not independent from $A$.
@@
We now use 4.3 to derive an algorithmic version of a result by Reed and Wood [26, Lemma 4.3].
@@
By the assumption that each $G[V_{i}\cup V_{j}]$ is $b$-degenerate and since every $b$-degenerate graph with $\ell$ vertices has at most $b\cdot\ell$ edges, we conclude that, for each $i\in[r]$, the number of edges with exactly one endpoint in $V_{i}$ is at most $(r-1)b\cdot 2t$. This, in turn, implies that $|\Gamma(A_{u,v})|\leq 4b(r-1)$ since, by definition, $A_{x,y}\in\Gamma(A_{u,v})$ if $\{x,y\}\cap\{u,v\}\neq\emptyset$ and $\{x,y\}\neq\{u,v\}$. The intuition here is that, for $u\in V_{i}$ and $v\in V_{j}$ with $i\neq j$, the event $A_{u,v}$ is independent from every other event in $\mathcal{A}$ associated with an edge $e\in F$ with endpoints in sets of $\mathcal{V}$ distinct from both $V_{i}$ and $V_{j}$, and thus the degeneracy assumption implies that the bad events are reasonably independent from each other. Defining $d=4b(r-1)t$, we are ready to invoke 4.3.
Since every $\textsf{Pr}(A_{u,v})=1/t^{2}$, to apply 4.3 it suffices to choose $x(A_{u,v})=1/d$ for all $(A_{u,v}\in\mathcal{A})$. Now, we want that
@@
and thus the choice of $t$ guarantees that the conditions of 4.3 are satisfied and a good evaluation is found in polynomial time.
@@ 4.3 The algorithmic construction
Let $G$ be a graph with one vertex for each linkage $L_{i,j}$ of the path system $\mathcal{(P,L,A)}$ with $i\neq j$ and no edges. Formally, $V=\{(i,j)\in[a]\times[a]\mid i\neq j\}$ and $G=(V,\emptyset)$. The authors of [20] apply [20, Lemma 3.5] to $\mathcal{(P,L,A)}$ and thus associate each vertex of $G$ with a threaded linkage of size $b$, as defined in [20]. The goal now is to apply [20, Lemma 3.6] to each of those objects to associate with each vertex of $G$ either a set of closed walks with additional properties, including small congestion, or with an untangled threaded linkage. For that, we need that $b\geq x\cdot d+(d-1)$, for some choice of $x$ and $d$, in order to conclude that the obtained sets of closed walks have size $d$ and the untangled threaded linkages have size at least $x$. Let $\alpha>1$. Choosing
@@
Thus, the lemma is applicable. The set of vertices of $G$ for which the first outcome is obtained, namely the closed walks, is named $Z$. As a consequence of applying [20, Lemma 3.6] to each $(i,j)\in Z$, we can also associate each such vertex with a threaded linkage satisfying some additional properties. Whether each $(i,j)\in Z$ is used to identify one of such threaded linkages or a closed walk depends on the application. For now,
@@
We remark that, so far, we have applied [20, Lemmas 3.5 and 3.6] a total of $a(a-1)$ times each.
They now build two undirected graphs on top of $V$, each depending on the choices of $d_{1}$ and $d_{2}$ with $d_{1}\geq d_{2}$. These choices are done in order to satisfy the LLL conditions, and thus need to be adapted to satisfy the polynomial LLL conditions. Start with $E_{1}=E_{2}=\emptyset$. For each pair $(i,j),(i^{\prime},j^{\prime})\in V$ and $i\in\{1,2\}$, add the pair to $E_{i}$ if ${\sf Int}(L^{\prime}_{i,j}\cup L^{\prime}_{i^{\prime},j^{\prime}})$ is not $d_{i}$-degenerate and define $H_{i}=(V,E_{i})$. Finally, let $M_{1}$ be a maximum matching in $H_{1}\setminus Z$ and $M_{2}$ be a maximum matching in $H_{2}$ excluding every edge with both endpoints in $Z$ plus vertices used by $M_{1}$. For $i\in\{1,2\}$ let $V(M_{i})$ be the set of vertices used by the edges of $M_{i}$. It is shown [20, Claim 4.1] that at least one of these three cases occurs:
@@
The algorithmic version of the sparse winning scenario is obtained by simply following the original proof and applying 4.1 instead of [20, cf. Lemma 2.4].
Lemma 4.6 (sparse winning scenario [20, Lemma 3.2], algorithmic version).
@@
Similarly, the algorithmic version of the sparse winning scenario (wrapped) is obtained by following the original proof and adjusting the bounds to apply Theorem 4.4 instead of [20, Lemma 2.5].
Lemma 4.7 (sparse winning scenario, wrapped [20, Lemma 3.3], algorithmic version).
Let $c$ be the constant from 4.6. Let $D$ be a digraph, let $a,b\geq 1$ be integers and $(P,\mathcal{L},\mathcal{A})$ be an $(a,b)$-path system in $D$. Let $\mathcal{I}$ be a subset of $[a]\times[a]\setminus\{(i,i)\mid i\in[a]\}$ with size at least $0.6a(a-1)$, and assume that there is an integer $d$ such that, for every distinct pair $(i,j),(i^{\prime},j^{\prime})\in\mathcal{I}$ the intersection graph ${\sf Int}(L_{i,j}\cup L_{i^{\prime},j^{\prime}})$ is $d$-degenerate. If $b^{1-\varepsilon}\geq(e\cdot 4a^{2}\cdot d)^{1+\varepsilon}$, then in polynomial time we can find in $D$ a bramble of congestion at most $4$ and size at least
@@
In addition, for case 2. another computation step is needed in the form of the Bowtie Lemma [20, Lemma 3.7], whose proof also naturally yields a polynomial-time algorithm, and the dense winning scenario. Again, the algorithmic version follows by incorporating 4.1 to the original proof.
Lemma 4.8 (dense winning scenario [20, Lemma 3.1], algorithmic version).
@@
For case 1., let $\mathcal{I}=V\setminus(V(M_{1})\cup Z)$. The goal is to apply 4.7 to a path system built from linkages associated with pairs $(i,j)\in\mathcal{I}$. Since the pairs in $Z$ are excluded, every linkage in $\{L^{\prime}_{i,j}\mid(i,j)\in\mathcal{I}\}$ has size at least $(e\cdot 4a^{2}\cdot d_{1})^{\alpha}+1$. Thus we need that